First cost of equipment = $150,000 Market value at the end of year 6 = $30,000 MACRS depreciation is used. The equipment is athe answer is no $43,000
Answer #1

Before tax cash flows for year 1 = BTCF = (Revenue - O&E expenses) = 60K = $60000

Since the equipment is a 5 year property, so Under MACRS

Depreciation for a year = Rate of depreciation x First cost of equipment

Rate of depreciation for year 1 under 5 year MACRS = 20%

Depreciation for year 1 = Rate of depreciation for year 1 under 5 year MACRS x First cost = 20% x 150000 = $30000

After tax cash flows for year 1 = (Revenue - O&M expenses - depreciation)(1-tax rate) + depreciation = (BTCF - Depreciation)(1-tax rate) + Depreciation = (60000 - 30000)(1-35%) + 30000 = 30000 x 65% + 30000 = 19500 + 30000 = $49500

So After tax cash flow for first year = $49500